# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ 3.1 Constraints
Cross-section: The interaction cross-section between the DM particles for this model is given by [38]
@@
The small-scale problems in the $\Lambda$CDM model and cold DM-only simulations like the i) cusp-core problem, ii) missing satellite problem, iii) too big to fail problem are known to be resolved by allowing for self-interacting DM (SIDM) [42]. These studies demand that the interaction strength be satisfied to a particular range in order to resolve the problems. The bounds on $\sigma/m$ are given by $1000.1<\sigma/m<$ 100 cm2/g. The upper limit is usually known to be around 10cm2/g, however we consider here least constraining case allowing for more flexibility and the largest possible parameter space. Thus,
@@
The upper limits come primarily from observations of merging clusters (for e.g., Bullet Cluster), whereas the lower bounds arise from the assumption that self-interactions alone are responsible for the resolution of the aforementioned problems encountered in CDM-only simulations (see [42] for a comprehensive list of all constraints). There are alternative solutions to these problems, the most obvious one being the inclusion of baryonic feedback at small scales. Thus, the lower bound is not strict, and hence, we denote it by a dashed line in Fig. 1. Only parameters consistent with these constraints were considered in [31]. However, we see a large range of parameter space is allowed, and stellar mass BSs can be produced even when the lower bound is removed.
@@
Seoane et al. [36] added constraints on the parameters space, assuming the object at the centre of our Milky Way is a BS (acting as a BH mimicker). This gave $3.4^{4}\leq m/\lambda^{4}\leq 2.9\times 10^{5}$ [eV]. A similar analysis for NGC 4258 gave $6.3\times 10^{4}\leq m/\lambda^{4}\leq 9.6\times 10^{5}$ [eV]. However, we do not make such an assumption while exploring the parameter space. Also, it is now established that Sagittarius A* is an SMBH [43], which is also true for the M87 galactic centre [44]. Arbey et al. [45] fitted scalar DM to galaxy rotation curves of dwarf spiral galaxies and obtained $7550<m^{4}/\lambda<$ 75 [eV4]. This corresponds to $M_{max}$ of $\mathcal{O}(10^{16})$, which is too large to be considered for our study of compact BSs. Eby et al. [38] impose the constraint of $\lambda\lesssim 4\pi$ for perturbative reasons, which, when combined with $\sigma/m$ constraint, allow DM of mass of only below 100 MeV approximately. We do not consider this condition here.
@@
Mass: Fuzzy DM can be described by ultra-light ALPs roughly in the mass range $m\in[10^{-25},10^{-17}]$ eV. Planck Collaboration put a lower bound on the DM mass of fuzzy DM as $m\gtrsim 10^{-25}$ eV using the CMB anisotropies. This has been updated to $m\gtrsim 10^{-23}$ eV by combining these results with Dark Energy Survey- Year 1 data[46] for weak gravitational lensing and to $m\gtrsim 2\times 10^{-20}$ eV using the Lyman-alpha forest [47]. More bounds were added from the study of the number of satellite galaxies for Milky Way as $m\gtrsim 2.9\times 10^{-21}$ eV [48] and $m\gtrsim 2.5\times 10^{-22}$ eV from the study of UV luminosities of high-redshift galaxies using Hubble and James-Webb telescopes [49]. These bounds on Bosonic DM assume axion potentials, which mimic massive BS potential for weak coupling.
@@
For the quartic potential case, [39] derived a lower bound of $m\gtrsim 10^{-24}$ eV using CMB and large-scale structure data of Planck Collaboration [50] and WiggleZ Dark Energy Survey [51]. There is a corresponding upper bound on $\lambda$ given as $\lambda/m^{4}<10^{20.14}$ [MeV-4]. This corresponds to $M_{max}\sim\mathcal{O}(10^{15})M_{\odot}$, which is much larger than the masses we are interested in for BSs. We still show it by a red line in Fig. 1. This also corresponds to an upper bound in $x$ given by $x<10^{10.07}$. These constraints were used to study the admixture of Bosonic DM in NS in [52].
The bounds discussed so far are model-dependent. Recently, Zimmerman et al. [40] derived a robust bound of $m\gtrsim 2.2\times 10^{-21}$ eV using dwarf galaxies, which does not depend on the assumed cosmology, microphysics, or dynamics of dark matter. The only assumption is that the DM is composed of a single scalar degree of freedom. This bound is shown in Fig. 1 as a blue-hatched exclusion region.
The recent discovery of stochastic GW background by the International Pulsar Timing Array (IPTA) also opened up a way to probe ultralight DM (ULDM). The array is sensitive to DM in the mass range $10^{-24}\lesssim m\lesssim 10^{-20}$ eV. The NANOGrav collaboration found no significant evidence for ULDM in the range $10^{-24}\lesssim m\lesssim 10^{-20}$ eV, while the EPTA collaboration concluded that ULDM in the mass range $10^{-24}\lesssim m\lesssim 10^{-22.3}$ eV cannot alone explain the DM budget. The bound by [40] is still higher than these, so we do not consider these.
@@ 4 Results
4.1 Scaled Relations - I (static observables)
We first discuss the scaling of the mass-radius relations along with compactness. This scaling property in the strong-interaction limit was first shown by Colpi et al. [12]. The self-similarity of the Tolman-Oppenheimer-Volkoff (TOV) equations was shown in [41]. In terms of the scaled quantities, the maximum mass observed is $M^{\prime}_{max}=0.06$, and the maximum compactness is $C_{max}=0.16$. For BSs, we can, in principle, go arbitrarily low in mass and compactness. In such a case, the radius reached a saturation value of $R^{\prime}_{max}=0.625$.
We derive analytical fit relations for the scaled $M^{\prime}-C$, $R^{\prime}-C$, and $M^{\prime}-R^{\prime}$ curves obtained by solving the scaled TOV equation (see Appendix of [41]), which we did not find in any previous literature. A recent work [53] provided fits for specific regions of the mass curve of mini-BSs (Kaup limit). Fit relations would enable switching between the quantities $M$, $R$, and $C$ using simple analytic expressions. Compactness is the most natural control variable for BSs as it does not scale with model parameter $x$. For this reason, we provide analytical fits for all quantities in terms of compactness wherever feasible.
In Fig.2(a), we plot the scaled mass $M^{\prime}$ as a function of compactness $C$. $M^{\prime}$ increases with $C$ from (0,0) and reaches a values of $M^{\prime}_{max}\approx 0.06$ at $C^{\prime}_{max}\approx 0.16$. We perform fitting only up to this critical point as the region after it, where the mass decreases, is unstable and not of interest. We fit quadratic and quartic functions of the form
@@ 4.1 Scaled Relations - I (static observables)
For this fit, we take $Y=M^{\prime}$ and $X=C$. The fit coefficients are given in Table. 1. We solve for $f$-modes for a minimum compactness of 0.02 in this work, so we restrict ourselves to $C\gtrsim 0.02$. This is lower than the minimum compactness of $C=0.035$ considered in previous works [54, 55]. For $C\gtrsim 0.02$, we observe that the quadratic functions provide a good fit with an accuracy within $2\%$. For the case of quartic fit, the accuracy is within $0.25\%$. Using the fit for $M^{\prime}-C$, we get $M^{\prime}(C=0.02)\approx 0.012$. So, in this work, we focus on BS configurations with $M^{\prime}\gtrsim 0.012$.
We also plot the relation given in Eq. A2 of [56] (see Eq. C.1) in yellow. This fit was derived for zero spin while studying the variation of $C^{-1}$ for different mass configurations as a function of dimensionless spin $\chi=J/M^{2}$, where $J$ is the spin angular momentum of BS. From the plot, we see that this does not fit the $M-C$ curve for non-rotating BS. This also results in a departure from expected values of $k_{2}$ (see Fig. 6 of [56]) as shown in Fig.14. Thus, our fit provides a new analytical relation for the same.
Next, we fit the scaled radius as a function of compactness, i.e., $Y=R^{\prime}$, $X=C$. We show the dependence in Fig. 2(b). As we increase the compactness, the radius decreases. The scaled radius decreases from $R^{\prime}=0.6$ for $C=0.02$ down to a minimum value of $R^{\prime}_{min}=0.38$. We show the quadratic and quartic fits as described in Eqn. 4.1 in blue and green, respectively. The accuracy is sub percent level, with that of the quadratic fit being under $1\%$ and that of the quartic fit being under $0.1\%$ for $C\gtrsim 0.02$. The fit coefficients are reported in Table. 1.
We show the well-known $M^{\prime}-R^{\prime}$ curve in Fig.2(c). Here $Y=M^{\prime}$, $X=R^{\prime}$. At low, masses ($M^{\prime}\to 0$), the radius reaches a constant values of $R^{\prime}_{max}=0.626$. $R^{\prime}$ decreases as $M^{\prime}$ increases and reaches a minimum value of $R^{\prime}=0.38$ as $M^{\prime}$ reaches a maximum value of $M^{\prime}=0.06$. We perform fits using the same functions as given in 4.1 with $y=M^{\prime}/M^{\prime}_{max}=M/M_{max}$ and $x=R^{\prime}/R^{\prime}_{max}=R/R_{max}$. The fit coefficients are given in Table. 1. So below $R^{\prime}(C=0.02)=0.6$, the quadratic function fits within an accuracy of $3.5\%$, and the quartic function within $0.75\%$.
The dimensionless tidal deformability ($\Lambda$) does not scale with $x$. This has been derived in Appendix A. We show how the quantities above $(M,R,C)$ vary with the central density in Appendix. B. We show the $\Lambda-M^{\prime}$ solution in Appendix. C and compare it with the previous results so far. The scaling for the moment of inertia has been derived in Appendix. D.
@@
Table 1: Fitting coefficients for the quadratic and quartic function fits to $M-C$, $M-R$, and $R-C$ as given in Eqs. 4.1. These coefficients are used to plot the fitting functions in Figs. 2 and 2.
@@
Figure 2: Quadratic (blue) and quartic fits (green) for (a) $M^{\prime}-C^{\prime}$, (b) $R^{\prime}-C^{\prime}$, and (c) $M^{\prime}-R^{\prime}$ relations for massive BSs in the strong-interaction limit. The dashed black line is the actual calculated numerical solution. The yellow curve is taken from [56], where the fit is derived for non-spinning massive BS. The percent errors for both fits are shown in the lower panels of each figure. The star marks the point beyond which the BS configurations become unstable. The stable solutions are valid only for compactness (radius) below (above) the star.
4.2 Scaled Relations - II (fundamental modes)
@@
We now move to $f$-modes, which is the focus of our work. We showed in [34] that upon applying the scaling $f^{\prime}=fxM_{Pl}$ and $\tau^{\prime}=\tau/xM_{Pl}$, we get unique solutions for $f$-mode characteristics as a function of $M^{\prime}$ and $C$. This is because the $f$-mode equations and perturbation equations are completely independent of model parameters when written in scaled coordinates. We have shown this explicitly in the Appendix of our other work [34]. We can use then scaling to obtain the $f-M$ curve for any set of $(\lambda,m)$ parameters.
@@
We show this curve as obtained in [34] in Fig. 3(a) with a black-dashed line. $f^{\prime}$ increases with $M^{\prime}$ and reaches a maximum value of $f^{\prime}_{max}=0.21$ for the maximum mass configuration of BS. We observe that the curve begins with a constant slope at low mass, and the slope diverges as it reaches the maximum mass. Hence, we fit a hyperbola to this relation given by
$f^{\prime}=-7.297\sqrt{\left(\frac{M^{\prime}-M^{\prime}_{max}-189.0116}{189.0116}\right)^{2}-1}+f^{\prime}_{max}~{}.$ (4.2)
@@
Here we have imposed the condition that $f^{\prime}(M^{\prime}=M_{max})=f^{\prime}_{max}$. The error of this fit is shown in the lower panel of Fig. 3(a). This fit agrees with the numerical solution within $3\%$. We also show the solution for $f^{\prime}-C^{\prime}$ in Fig. 3(b) as compactness is a more natural variable. This, too, is an increasing function, and we fit the quadratic and quartic functions given in Eqn. 4.1. Here $Y=f^{\prime}$ and $X=C$. The quadratic fit diverges for both low and high compactness. The quartic fit is accurate to $1\%$ for $C\gtrsim 0.02$. The fit coefficients are reported in Table. 2.
@@
Figure 3: The scaled dimensionless $f$-mode frequency ($f^{\prime}$) as a function of a) scaled mass of BSs and b) compactness. The blue curve in the left panel is the fitted hyperbola as given in Eq. 4.2. The blue and the green curves in the right panel are the fits for the quadratic and quartic functions, respectively, as given in Eqn. 4.1. The fit coefficients are given in Table. 2. The lower panel shows the percent error for each fit.
Next, we show $\tau^{\prime}$ as a function of $M^{\prime}$ and $C^{\prime}$ as reported in [34] in Fig. 4. $\tau^{\prime}$ is decreasing function of $M^{\prime}$ and $C$ as expected. We use the log scale for $\tau^{\prime}$ as it varies by orders of magnitude. The minimum value of $\tau^{\prime}$ corresponding to the maximum mass configuration is $\tau^{\prime}_{min}=1900$ [34]. This single fit is sufficient to reproduce all the results and for the entire parameter space in the strong coupling regime ($\Lambda\gg 1$) using scaling relations.
For both cases, we fit the quartic polynomial of the form Eq. 4.1. For the $\tau^{\prime}-M^{\prime}$ fit we use $X=1/(\log{\tau^{\prime}})$ and $Y=M^{\prime}$ as in the case of $\Lambda-M^{\prime}$ relation [57, 56]. We find the error is within $5\%$. For the case of $\tau^{\prime}-C^{\prime}$, choose $X=1/C^{\prime}$ and $Y=\tau^{\prime}$. The fit has a percent level of accuracy. The fit coefficients are provided in Table. 2.
Figure 4: Same as Fig. 3 but for the scaled dimensionless $f$-mode damping time ($\tau^{\prime}$).
@@
$f^{\prime}$ $C^{\prime}$ 3.747 ×$\times 10^{-2}$ 0.656 2.627 2.363 ×$\times 10^{-2}$ 1.203 -2.232 -4.417 113.160
$M^{\prime}$ $1/\log{\tau^{\prime}}$- -- -- -9.718 ×$\times 10^{-2}$ 4.457 -84.357 7.771 ×$\times 10^{2}$2.460 ×$\times 10^{3}$
$\tau^{\prime}$ $1/C^{\prime}$- -- -- 4.124 ×$\times 10^{2}$1.147 ×$\times 10^{2}$ 24.602 5.030 3.856 ×$\times 10^{-3}$

### v3

@@ 3.1 Constraints
There are various independent constraints on the $\lambda$ and $m$ parameters from various astrophysical observations. We provide a brief review of these constraints below.
Cross-section: The self-interaction cross-section ($\sigma$) between the DM particles is related to the $\lambda$ parameter of this model by [Eby2016]
@@
The upper limits come primarily from observations of merging clusters (for e.g., Bullet Cluster), whereas the lower bounds arise from the assumption that self-interactions alone are responsible for the resolution of the aforementioned problems encountered in CDM-only simulations (see [Tulin2018] for a comprehensive list of all constraints). There are alternative solutions to these problems, the most obvious one being the inclusion of baryonic feedback at small scales. Thus, the lower bound is not strict, and hence, we denote it by a dashed line in Fig. 1. Only parameters consistent with these constraints were considered in [Flores2019]. However, we see that a large range of parameter space is allowed, and stellar mass BSs can be produced even when the lower bound is removed.
@@
Seoane et al. [Seoane2010] added constraints on the parameter space, assuming the object at the centre of our Milky Way is a BS (acting as a BH mimicker). This gave $3.7\times 10^{4}\leq m/\lambda^{1/4}\leq 2.9\times 10^{5}$ [eV]. A similar analysis for NGC 4258 gave $6.3\leq m/\lambda^{1/4}\leq 9.6\times 10^{4}$ [eV]. However, we do not make such an assumption while exploring the parameter space. Also, it is now established that Sagittarius A* is a supermassive BH (SMBH) [SagittariusEHT2022], which is also true for the M87 galactic centre [M87EHT2019]. Arbey et al. [Arbey2003] fitted scalar DM to galaxy rotation curves of dwarf spiral galaxies and obtained $50<m^{4}/\lambda<75$ [eV4]. This corresponds to $M_{\rm max}$ of $\mathcal{O}(10^{16})M_{\odot}$, which is too large to be considered for our study of compact BSs. Eby et al. [Eby2016] impose the constraint of $\lambda\lesssim 4\pi$ for perturbative reasons, which, when combined with $\sigma/m$ constraint, allow DM of mass of only below 100 MeV approximately. We do not consider this condition here.
@@
Mass: Fuzzy DM can be described by ultra-light ALPs roughly in the mass range $m\in[10^{-25},10^{-17}]$ eV. Planck Collaboration put a lower bound on the DM mass of fuzzy DM as $m\gtrsim 10^{-24}$ eV using the CMB anisotropies [Hlozek2015]. This has been updated to $m\gtrsim 10^{-23}$ eV by combining these results with Dark Energy Survey- Year 1 data[Dentler2022] for weak gravitational lensing and to $m\gtrsim 2\times 10^{-20}$ eV using the Lyman-alpha forest [Rogers2021]. More bounds were added from the study of the number of satellite galaxies for Milky Way as $m\gtrsim 2.9\times 10^{-21}$ eV [Nadler2021] and $m\gtrsim 2.5\times 10^{-22}$ eV from the study of UV luminosities of high-redshift galaxies using Hubble and James-Webb telescopes [Winch2024]. These bounds on Bosonic DM assume axion potentials, which mimic a massive BS potential for weak coupling.
@@
For the quartic potential case, [Cembranos2018] derived a lower bound of $m\gtrsim 10^{-24}$ eV using CMB and large-scale structure data of Planck Collaboration [PlanckCollaboration2016] and WiggleZ Dark Energy Survey [Parkinson2012]. There is a corresponding upper bound on $\lambda$ given as $\lambda/m^{4}<10^{20.14}$ [MeV-4]. This corresponds to $M_{\rm max}\sim\mathcal{O}(10^{15})M_{\odot}$, which is much larger than the masses we are interested in for BSs. We still show these by a red line in Fig. 1. This also corresponds to an upper bound in $x$ given by $x<10^{10.07}$. Since we consider a quartic potential here, this is the most relevant constraint. These constraints were used to study the admixture of Bosonic DM in NSs in [Rezaei2023].
The bounds discussed so far are model-dependent. Many independent bounds for DM come from studies of fuzzy DM that consider non-interacting scalar particles. Various constraints on DM mass come from observations of dwarf Milky Way satellite galaxies [Hayashi2021, Zimmerman2024, Teodori2025, Benito2025], their number [Nadler2021], dynamical heating [Marsh2019, Dalal2022], stellar streams [Benito2020], lensed radio jet [Powell2023], lensed quasars [Laroche2022], 21-cm observations [Nebrin2019], high redshift galaxies [Kulkarni2022], SMBH superradiance [Stott2018, Davoudiasl2019] and the recently discovered stochastic GWs [Afzal2023Nanograv, Smarra2023] using the International Pulsar Timing Array (IPTA). Each of these methods involves different assumptions and systematics, providing independent and complementary constraints on DM mass. Out of these, Zimmerman et al. [Zimmerman2024] is a robust bound of $m\gtrsim 2.2\times 10^{-21}$ eV using dwarf galaxies, which does not depend on the assumed cosmology, microphysics, or dynamics of dark matter. The only assumption is that the DM is composed of a single scalar degree of freedom. This bound was recently improved further to $m\gtrsim 5\times 10^{-21}$ eV by Teodori et al. [Teodori2025], taking into account the dynamics as opposed to using only kinematic information as in the case of the former. However, the former still remains model-independent and robust. The strongest constraints are obtained from the study of dynamical heating of stellar orbits in ultra-faint dwarf galaxies [Dalal2022], which imposes $m>3\times 10^{-19}$ eV. Note that all these bounds do not consider any self-interactions. We consider the most robust and model-independent one here [Zimmerman2024] only for the purpose of depiction. This bound is shown in Fig. 1 as a blue-hatched exclusion region.
@@ 4 Results
In this section, we first present the solution for the scaled static variables $M^{\prime}$, $R^{\prime}$, and $C$. We then focus on the $f$-mode characteristics and their universal relations with various macroscopic observables. Finally, we study the DM parameter space that is sensitive to various gravitational wave detectors and discuss the prospects of detectability.
4.1 Scaled Relations for static observables
We first discuss the scaling of the mass-radius relations along with compactness. Although these have been studied numerically before, we provide analytical fits to these curves, which we did not find in any previous literature. Fit relations would enable switching between the quantities $M$, $R$, and $C$ using simple analytic expressions and can be directly used in future studies on massive BSs. These solutions are obtained by considering the scaled EoS (see Eq. 2.5) and solving the scaling property of the TOV equation (see appendix of [Maselli2017, Celato2025] in the strong-interaction limit, which was first shown by Colpi et al. [Colpi1986].
In Fig.2(a), we plot the scaled mass $M^{\prime}$ as a function of compactness $C$. $M^{\prime}$ increases with $C$ and reaches a value of $M^{\prime}_{\rm max}\approx 0.06$ at $C^{\prime}_{\rm max}=C_{\rm max}\approx 0.16$. We perform fitting only up to this critical point, as the region after it, where the mass decreases, is unstable and not of interest. We fit quadratic and quartic functions of the form
@@ 4.1 Scaled Relations for static observables
The fit coefficients are given in Table. 1. We solve for $f$-modes for a minimum compactness of 0.02 in this work, so we restrict ourselves to $C\gtrsim 0.02$. This is lower than the minimum compactness of $C=0.035$ considered in previous works [Ryan1997, Vaglio2023]. For $C\gtrsim 0.02$, we observe that the quadratic functions provide a good fit with an accuracy within $2\%$. For the case of the quartic fit, the accuracy is within $0.25\%$. Note that these are fitting errors signifying the deviation of the analytical fit from the numerical solution.
We also plot the relation given in Eq. A2 of [Pacilio2020] in yellow. This fit was derived for zero spin while studying the variation of $C^{-1}$ for different mass configurations as a function of dimensionless spin $\chi=J/M^{2}$, where $J$ is the spin angular momentum of BS. From the plot, we see that this does not fit the MCM-C curve for a non-rotating BS. This also results in an unexpected behaviour of $k_{2}$ (Fig. 6 of [Pacilio2020]). A possible reason could be that the fit in [Pacilio2020] was obtained using only five configurations in the low spin limit. Thus, our fit provides a new analytical relation for the same.
We show the well-known $M^{\prime}-R^{\prime}$ curve in Fig.2(b). At low masses, the radius reaches a constant value of $^{\prime}_{\rm max}=0.626. RR^{\prime}$ decreases and reaches a minimum value of $R^{\prime}=0.38$ as $M^{\prime}$ increases reaching a maximum value of $M^{\prime}=0.06$. Below $R^{\prime}(C=0.02)=0.6$, the quadratic function fits within an accuracy of $3.5\%$, and the quartic function within $0.75\%$. We also perform the fit for $R^{\prime}-C$ (not shown here). The fit coefficients for both these relations are provided in Table. 1.
The dimensionless tidal deformability ($\bar{\Lambda}$) does not scale with $x$. This has been derived in Appendix A. The scaling for the moment of inertia has been derived in Appendix. B.
@@
Table 1: Fitting coefficients for the quadratic and quartic function fits to $M^{\prime}-C^{\prime}$, $M^{\prime}-R^{\prime}$, and $R^{\prime}-C^{\prime}$ as given in Eqs. 4.1. These coefficients are used to plot the fitting functions in Figs. 2 and 2.
@@
Figure 2: Quadratic (blue) and quartic fits (green) as described in Eqn. 4.1 for (a) $M^{\prime}-C^{\prime}$, and (b) $M^{\prime}-R^{\prime}$ relations for massive BSs in the strong-interaction limit. The dashed black line is the actual calculated numerical solution. The yellow curve is taken from [Pacilio2020], where the fit is derived for non-spinning massive BSs. The percent errors for both fits are shown in the lower panels of each figure. The star marks the point beyond which the BS configurations become unstable. The stable solutions are valid only for compactness (radius) below (above) the star. The fit coefficients are reported in Table. 1.
4.2 Scaled Relations for fundamental modes
@@
We now move to $f$-mode characteristics, which are the focus of our work. We showed in [shirke_BS_fmodes_short] that upon applying the scaling $f^{\prime}=fxM_{\rm Pl}$ and $\tau^{\prime}=\tau/xM_{\rm Pl}$, we get unique solutions for $f$-mode characteristics as a function of $M^{\prime}$ and $C$. This is because the $f$-mode equations and perturbation equations are completely independent of microscopic model parameters when written in scaled coordinates. We have shown this explicitly in the appendix of [shirke_BS_fmodes_short]. Just like the TOV solutions, for every central density $\rho^{\prime}$, we get a value for $f^{\prime}$ and $\tau^{\prime}$ and obtain a curve as we vary the central density. Here, we again provide analytical fits to the solutions numerically obtained in [shirke_BS_fmodes_short] for easier future computations. We can then use scaling to obtain the fMf-M curve for any set of $(\lambda,m)$ parameters.
@@
We show this curve as obtained numerically in [shirke_BS_fmodes_short] in Fig. 3(a) with a black-dashed line. This is obtained by solving for $f$-mode characteristics as discussed in Sec. 2. $f^{\prime}$ increases with $M^{\prime}$ and reaches a maximum value of $f^{\prime}_{\rm max}=0.21$ for the maximum mass configuration of the BS. We observe that the curve begins with a constant slope at low mass, and the slope diverges as it reaches the maximum mass. Hence, we fit a hyperbola to this relation given by
$f^{\prime}=-7.297\sqrt{\left(\frac{M^{\prime}-M^{\prime}_{\rm max}-189.0116}{189.0116}\right)^{2}-1}+f^{\prime}_{\rm max}~.$ (4.2)
@@
Here we have imposed the condition that $f^{\prime}(M^{\prime}=M_{\rm max})=f^{\prime}_{\rm max}$. This analytical fit agrees with the numerical solution within $3\%$. We also show the solution for $f^{\prime}-C^{\prime}$ in Fig. 3(b) as compactness is a more natural variable. The quartic fit is accurate to $1\%$ for $C\gtrsim 0.02$.
@@
Figure 3: The scaled dimensionless $f$-mode frequency ($f^{\prime}$) as a function of a) scaled mass and b) compactness of BSs. The blue curve in the left panel is the fitted hyperbola as given in Eq. 4.2. The blue and the green curves in the right panel are the fits for the quadratic and quartic functions, respectively, as given in Eqn. 4.1. The fit coefficients are given in Table. 2. The lower panel shows the percent error for each fit.
Similar fits can be performed for the damping time (figures not shown). The fit coefficients are provided in Table. 2.
@@
$f^{\prime}$ $C^{\prime}$ 3.747 $\times 10^{-2}$ 0.656 2.627 2.363 $\times 10^{-2}$ 1.203 -2.232 -4.417 113.160
$M^{\prime}$ $1/\log{\tau^{\prime}}$ - - - -9.718 $\times 10^{-2}$ 4.457 -84.357 7.771 $\times 10^{2}$ -2.460 $\times 10^{3}$
$\tau^{\prime}$ $1/C^{\prime}$ - - - 4.124 $\times 10^{2}$ -1.147 $\times 10^{2}$ 24.602 5.030 3.856 $\times 10^{-3}$